# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

### v1

@@ Appendix A Appendix
A.1 Code
The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
@@ A.2 SIB-Fleurs
Language Train Validation Test afr_Latn 406 86 95 amh_Ethi 752 54 149 arb_Arab 579 64 133 asm_Beng 730 71 176 ast_Latn 701 69 177 azj_Latn 712 71 174 bel_Cyrl 690 71 177 bul_Cyrl 749 70 176 ben_Beng 742 71 176 bos_Latn 746 71 177 cat_Latn 683 71 177 ceb_Latn 741 61 149 ckb_Arab 738 70 176 zho_Hans 751 71 176 ces_Latn 732 68 172 cym_Latn 739 71 177 dan_Latn 696 70 177 deu_Latn 736 69 175 ell_Grek 750 67 168 eng_Latn 738 71 177 spa_Latn 676 71 177 est_Latn 700 71 176 pes_Arab 692 66 165 fin_Latn 735 71 175 tgl_Latn 604 71 176 fra_Latn 753 65 164 gle_Latn 731 71 176 glg_Latn 660 71 174 guj_Gujr 752 71 177 hau_Latn 753 70 166 heb_Hebr 754 70 175 hin_Deva 653 60 132 hrv_Latn 756 71 176 hun_Latn 750 71 177 hye_Armn 741 71 177 ind_Latn 728 69 167 ibo_Latn 737 71 177 isl_Latn 381 18 23 ita_Latn 743 69 175 jpn_Jpan 662 62 164 jav_Latn 740 67 171 kat_Geor 557 69 177 kam_Latn 752 69 179 kea_Latn 725 71 175 kaz_Cyrl 749 70 176 khm_Khmr 588 69 168 kan_Knda 660 70 174 kor_Hang 669 61 141 kir_Cyrl 729 71 177 ltz_Latn 703 71 176 lug_Latn 691 70 173
Language Train Validation Test lin_Latn 755 59 139 lao_Laoo 591 54 132 lit_Latn 730 71 178 luo_Latn 698 39 98 lvs_Latn 634 69 174 mri_Latn 749 71 176 mkd_Cyrl 680 71 177 mal_Mlym 723 68 174 khk_Cyrl 743 71 177 mar_Deva 749 71 177 zsm_Latn 713 67 171 mlt_Latn 731 71 176 mya_Mymr 746 71 175 nob_Latn 723 51 127 npi_Deva 754 70 175 nld_Latn 729 58 123 nso_Latn 633 70 169 nya_Latn 720 68 169 oci_Latn 756 71 177 ory_Orya 442 71 168 pan_Guru 580 56 143 pol_Latn 723 68 165 pbt_Arab 701 55 144 por_Latn 728 70 177 ron_Latn 734 69 177 rus_Cyrl 733 71 173 snd_Arab 749 71 177 slk_Latn 628 71 169 slv_Latn 704 71 174 sna_Latn 689 71 176 som_Latn 746 70 177 srp_Cyrl 730 63 164 swe_Latn 686 71 168 swh_Latn 745 65 154 tam_Taml 693 71 169 tel_Telu 658 66 153 tgk_Cyrl 680 69 163 tha_Thai 710 71 176 tur_Latn 692 67 164 ukr_Cyrl 732 67 164 umb_Latn 473 39 108 urd_Arab 636 65 120 uzn_Latn 734 69 175 vie_Latn 737 70 176 wol_Latn 643 52 123 xho_Latn 756 71 177 yor_Latn 686 71 172 zho_Hant 624 70 172 zul_Latn 739 69 175 fuv_Latn 752 68 166 gaz_Latn 574 6 17
@@ A.3 Belebele-Fleurs
Language Samples afr_Latn 309 amh_Ethi 782 arb_Arab 387 asm_Beng 824 azj_Latn 759 ben_Beng 855 bul_Cyrl 873 cat_Latn 652 ceb_Latn 783 ces_Latn 802 ckb_Arab 842 dan_Latn 696 deu_Latn 804 ell_Grek 837 eng_Latn 844 est_Latn 736 fin_Latn 826 fra_Latn 839 fuv_Latn 848 gaz_Latn 25 guj_Gujr 880 hau_Latn 838 heb_Hebr 878 hin_Deva 515 hrv_Latn 896 hun_Latn 879 hye_Armn 861 ibo_Latn 838 ind_Latn 783 isl_Latn 81 ita_Latn 851 jav_Latn 835 jpn_Jpan 590 kan_Knda 606 kat_Geor 372 kaz_Cyrl 870 kea_Latn 770 khk_Cyrl 869 khm_Khmr 439 kir_Cyrl 811 kor_Hang 535 lao_Laoo 346 lin_Latn 778 lit_Latn 834 lug_Latn 703 luo_Latn 512
Language Samples lvs_Latn 555 mal_Mlym 809 mar_Deva 869 mkd_Cyrl 667 mlt_Latn 816 mri_Latn 877 mya_Mymr 864 nld_Latn 674 nob_Latn 635 npi_Deva 876 nso_Latn 569 nya_Latn 752 ory_Orya 220 pan_Guru 396 pbt_Arab 628 pes_Arab 673 pol_Latn 765 por_Latn 791 ron_Latn 815 rus_Cyrl 819 slk_Latn 513 slv_Latn 724 sna_Latn 735 snd_Arab 878 som_Latn 874 spa_Latn 659 srp_Cyrl 766 swe_Latn 681 swh_Latn 780 tam_Taml 714 tel_Telu 567 tgk_Cyrl 632 tgl_Latn 505 tha_Thai 745 tur_Latn 706 ukr_Cyrl 773 urd_Arab 482 uzn_Latn 812 vie_Latn 847 wol_Latn 495 xho_Latn 900 yor_Latn 652 zho_Hans 888 zho_Hant 527 zsm_Latn 749 zul_Latn 838
@@ A.5 Analysis of ASR Performance
Language Whisper-v3-Large SeamlessM4Tv2-Large afr_Latn 13.1 12.6 amh_Ethi 207.8 23.9 arb_Arab 7.6 8.0 asm_Beng 99.0 15.1 ast_Latn 15.5 19.8 azj_Latn 7.2 8.3 bel_Cyrl 11.9 6.9 ben_Beng 33.9 9.0 bos_Latn 4.6 7.8 bul_Cyrl 5.2 9.6 cat_Latn 3.3 5.3 ceb_Latn 17.4 19.0 ces_Latn 4.2 8.0 ckb_Arab 98.3 13.9 cym_Latn 15.8 14.4 dan_Latn 4.8 10.0 deu_Latn 2.2 6.5 ell_Grek 6.4 10.8 eng_Latn 3.1 6.2 est_Latn 7.3 11.0 fin_Latn 3.2 10.2 fra_Latn 2.8 6.6 fuv_Latn 194.3 57.0 gaz_Latn 35.8 47.8 gle_Latn 85.1 21.4 glg_Latn 4.3 6.5 guj_Gujr 21.3 10.3 hau_Latn 33.8 51.8 heb_Hebr 12.2 14.5 hin_Deva 11.7 10.2 hrv_Latn 4.5 8.5 hun_Latn 5.2 8.5 hye_Armn 18.0 9.1 ibo_Latn 42.9 59.3 ind_Latn 4.0 9.2 isl_Latn 16.0 10.3 ita_Latn 2.1 5.2 jav_Latn 26.3 11.3 jpn_Jpan 6.3 16.4 kam_Latn 45.1 53.9 kan_Knda 20.1 10.9 kat_Geor 19.8 7.1 kaz_Cyrl 8.9 9.7 kea_Latn 35.8 38.1 khk_Cyrl 37.6 14.6 khm_Khmr 144.0 29.8 kir_Cyrl 28.3 8.8 kor_Hang 7.4 10.3 lao_Laoo 109.7 30.3 lin_Latn 22.0 58.7 lit_Latn 8.3 11.9 ltz_Latn 29.5 40.1
Language Whisper-v3-Large SeamlessM4Tv2-Large lug_Latn 44.9 14.1 luo_Latn 42.7 58.8 lvs_Latn 7.3 8.2 mal_Mlym 104.3 14.2 mar_Deva 24.3 10.7 mkd_Cyrl 5.8 8.6 mlt_Latn 26.0 12.1 mri_Latn 13.4 54.7 mya_Mymr 132.9 22.2 nld_Latn 3.1 6.8 nob_Latn 4.6 9.0 npi_Deva 25.6 74.1 nso_Latn 95.7 66.9 nya_Latn 35.6 12.4 oci_Latn 25.9 33.8 ory_Orya 93.8 11.4 pan_Guru 44.2 10.3 pbt_Arab 36.8 21.6 pes_Arab 8.6 7.7 pol_Latn 2.3 7.9 por_Latn 2.8 7.6 ron_Latn 3.2 6.6 rus_Cyrl 2.6 6.1 slk_Latn 3.9 7.0 slv_Latn 6.0 9.2 sna_Latn 26.7 32.7 snd_Arab 94.5 33.5 som_Latn 34.9 17.9 spa_Latn 2.2 6.3 srp_Cyrl 76.6 7.6 swe_Latn 3.7 9.2 swh_Latn 11.0 8.9 tam_Taml 13.3 11.6 tel_Telu 90.1 13.3 tgk_Cyrl 28.8 8.8 tgl_Latn 5.0 12.0 tha_Thai 9.8 11.4 tur_Latn 6.2 7.5 ukr_Cyrl 3.0 8.9 umb_Latn 156.7 55.5 urd_Arab 9.4 9.6 uzn_Latn 25.0 8.4 vie_Latn 4.7 6.7 wol_Latn 140.6 47.9 xho_Latn 48.1 32.3 yor_Latn 49.3 33.2 zho_Hans 16.2 19.0 zho_Hant 30.6 35.1 zsm_Latn 3.4 8.5 zul_Latn 42.6 18.4

### v2

@@ 6 Conclusion
Limitations
Our work has two primary limitations. First, the selection of models in CS can be further expanded. In particular, additional evaluations on CS that combine transcription models with prompting 70B LLMs would provide valuable insights given the growing interest in instruction fine-tuned LLMs. Moreover, assessing speech-LLMs on both tasks would further illuminate multimodal alignment in massively multilingual scenarios. Notably, LLM2Vec on Llama 3.1 8B, which underwent extensive task fine-tuning on 63K instances, represents an exceptionally strong language model backbone for CS. Compared to concurrent work, our models would have even outperformed prompting 70B LLMs on speech transcriptions in Belebele Costa-jussà et al. (2024). Additionally, current speech LLMs rely on Whisper, which is inherently limited to 30-second speech segments. This limitation hinders their application in Belebele-Fleurs, where the average speech length is 47 seconds (cf. Table 2). Our compute infrastructure was unable to support prompting such large LLMs on long contexts, exceeding half a minute of speech combined with corresponding questions and answers. We therefore leave these evaluations for future work.
Furthermore, although our findings indicate a clear link between robust multilingual SLU and strong multilingual ASR, future research must more carefully isolate the mechanisms underlying this interaction. Such investigations should clarify the conditions (e.g., training objectives, data and language mixtures) under which SLU most effectively contributes to more robust multilingual ASR. Nonetheless, our analysis strongly suggests that multilingual SLU and multilingual ASR are mutually beneficial.
Acknowledgments
We use AI assistance, specifically chatGPT o3-mini, to refine both the writing and the tables in our manuscript.
@@ Appendix A Appendix
A.1 Further Experimental Details
Compute infrastructure. All experiments run on a single Nvidia L40S 48GB or A100 80GB, respectively. We estimate that the total compute budget accumulates to about 2K GPU hours. Training runs on SIB-Fleurs require about 30-45 minutes, while fine-tuning LLMs for Belebele require roughly 9 hours per run. Evaluation, per checkpoint, across all languages supported by the corresponding benchmark, requires about 1 hour due to the comprehensive setups (e.g., original text and two types of transcriptions for speech recordings).
Code. The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
@@ A.2 SIB-Fleurs
Language Code Language Name Train Validation Test afr_Latn Afrikaans 406 86 95 amh_Ethi Amharic 752 54 149 arb_Arab Modern Standard 579 64 133 asm_Beng Assamese 730 71 176 ast_Latn Asturian 701 69 177 azj_Latn North Azerbaijani 712 71 174 bel_Cyrl Belarusian 690 71 177 ben_Beng Bengali 742 71 176 bos_Latn Bosnian 746 71 177 bul_Cyrl Bulgarian 749 70 176 cat_Latn Catalan 683 71 177 ceb_Latn Cebuano 741 61 149 ces_Latn Czech 732 68 172 ckb_Arab Central Kurdish 738 70 176 cym_Latn Welsh 739 71 177 dan_Latn Danish 696 70 177 deu_Latn German 736 69 175 ell_Grek Greek 750 67 168 eng_Latn English 738 71 177 est_Latn Estonian 700 71 176 fin_Latn Finnish 735 71 175 fra_Latn French 753 65 164 fuv_Latn Nigerian Fulfulde 752 68 166 gaz_Latn West Central Oromo 574 6 17 gle_Latn Irish 731 71 176 glg_Latn Galician 660 71 174 guj_Gujr Gujarati 752 71 177 hau_Latn Hausa 753 70 166 heb_Hebr Hebrew 754 70 175 hin_Deva Hindi 653 60 132 hrv_Latn Croatian 756 71 176 hun_Latn Hungarian 750 71 177 hye_Armn Armenian 741 71 177 ibo_Latn Igbo 737 71 177 ind_Latn Indonesian 728 69 167 isl_Latn Icelandic 381 18 23 ita_Latn Italian 743 69 175 jav_Latn Javanese 740 67 171 jpn_Jpan Japanese 662 62 164 kam_Latn Kamba 752 69 179 kan_Knda Kannada 660 70 174 kat_Geor Georgian 557 69 177 kaz_Cyrl Kazakh 749 70 176 kea_Latn Kabuverdianu 725 71 175 khk_Cyrl Halh 743 71 177 khm_Khmr Khmer 588 69 168 kir_Cyrl Kyrgyz 729 71 177 kor_Hang Korean 669 61 141 lao_Laoo Lao 591 54 132 lin_Latn Lingala 755 59 139 lit_Latn Lithuanian 730 71 178
Language Code Language Name Train Validation Test ltz_Latn Luxembourgish 703 71 176 lug_Latn Ganda 691 70 173 luo_Latn Luo 698 39 98 lvs_Latn Standard Latvian 634 69 174 mal_Mlym Malayalam 723 68 174 mar_Deva Marathi 749 71 177 mkd_Cyrl Macedonian 680 71 177 mlt_Latn Maltese 731 71 176 mri_Latn Maori 749 71 176 mya_Mymr Burmese 746 71 175 nld_Latn Dutch 729 58 123 nob_Latn Norwegian Bokmål 723 51 127 npi_Deva Nepali 754 70 175 nso_Latn Northern Sotho 633 70 169 nya_Latn Nyanja 720 68 169 oci_Latn Occitan 756 71 177 ory_Orya Odia 442 71 168 pan_Guru Eastern Panjabi 580 56 143 pbt_Arab Southern Pashto 701 55 144 pes_Arab Western Persian 692 66 165 pol_Latn Polish 723 68 165 por_Latn Portuguese 728 70 177 ron_Latn Romanian 734 69 177 rus_Cyrl Russian 733 71 173 slk_Latn Slovak 628 71 169 slv_Latn Slovenian 704 71 174 sna_Latn Shona 689 71 176 snd_Arab Sindhi 749 71 177 som_Latn Somali 746 70 177 spa_Latn Spanish 676 71 177 srp_Cyrl Serbian 730 63 164 swe_Latn Swedish 686 71 168 swh_Latn Swahili 745 65 154 tam_Taml Tamil 693 71 169 tel_Telu Telugu 658 66 153 tgk_Cyrl Tajik 680 69 163 tgl_Latn Tagalog 604 71 176 tha_Thai Thai 710 71 176 tur_Latn Turkish 692 67 164 ukr_Cyrl Ukrainian 732 67 164 umb_Latn Umbundu 473 39 108 urd_Arab Urdu 636 65 120 uzn_Latn Northern Uzbek 734 69 175 vie_Latn Vietnamese 737 70 176 wol_Latn Wolof 643 52 123 xho_Latn Xhosa 756 71 177 yor_Latn Yoruba 686 71 172 zho_Hans Chinese 751 71 176 zho_Hant Chinese 624 70 172 zsm_Latn Standard Malay 713 67 171 zul_Latn Zulu 739 69 175
@@ A.3 Belebele-Fleurs
Language Code Language Name Samples afr_Latn Afrikaans 309 amh_Ethi Amharic 782 arb_Arab Modern Standard 387 asm_Beng Assamese 824 azj_Latn North Azerbaijani 759 ben_Beng Bengali 855 bul_Cyrl Bulgarian 873 cat_Latn Catalan 652 ceb_Latn Cebuano 783 ces_Latn Czech 802 ckb_Arab Central Kurdish 842 dan_Latn Danish 696 deu_Latn German 804 ell_Grek Greek 837 eng_Latn English 844 est_Latn Estonian 736 fin_Latn Finnish 826 fra_Latn French 839 fuv_Latn Nigerian Fulfulde 848 gaz_Latn West Central Oromo 25 guj_Gujr Gujarati 880 hau_Latn Hausa 838 heb_Hebr Hebrew 878 hin_Deva Hindi 515 hrv_Latn Croatian 896 hun_Latn Hungarian 879 hye_Armn Armenian 861 ibo_Latn Igbo 838 ind_Latn Indonesian 783 isl_Latn Icelandic 81 ita_Latn Italian 851 jav_Latn Javanese 835 jpn_Jpan Japanese 590 kan_Knda Kannada 606 kat_Geor Georgian 372 kaz_Cyrl Kazakh 870 kea_Latn Kabuverdianu 770 khk_Cyrl Halh 869 khm_Khmr Khmer 439 kir_Cyrl Kyrgyz 811 kor_Hang Korean 535 lao_Laoo Lao 346 lin_Latn Lingala 778 lit_Latn Lithuanian 834 lug_Latn Ganda 703 luo_Latn Luo 512 lvs_Latn Standard Latvian 555 mal_Mlym Malayalam 809
Language Name Samples mar_Deva Marathi 869 mkd_Cyrl Macedonian 667 mlt_Latn Maltese 816 mri_Latn Maori 877 mya_Mymr Burmese 864 nld_Latn Dutch 674 nob_Latn Norwegian Bokmål 635 npi_Deva Nepali 876 nso_Latn Northern Sotho 569 nya_Latn Nyanja 752 ory_Orya Odia 220 pan_Guru Eastern Panjabi 396 pbt_Arab Southern Pashto 628 pes_Arab Western Persian 673 pol_Latn Polish 765 por_Latn Portuguese 791 ron_Latn Romanian 815 rus_Cyrl Russian 819 slk_Latn Slovak 513 slv_Latn Slovenian 724 sna_Latn Shona 735 snd_Arab Sindhi 878 som_Latn Somali 874 spa_Latn Spanish 659 srp_Cyrl Serbian 766 swe_Latn Swedish 681 swh_Latn Swahili 780 tam_Taml Tamil 714 tel_Telu Telugu 567 tgk_Cyrl Tajik 632 tgl_Latn Tagalog 505 tha_Thai Thai 745 tur_Latn Turkish 706 ukr_Cyrl Ukrainian 773 urd_Arab Urdu 482 uzn_Latn Northern Uzbek 812 vie_Latn Vietnamese 847 wol_Latn Wolof 495 xho_Latn Xhosa 900 yor_Latn Yoruba 652 zho_Hans Chinese 888 zho_Hant Chinese 527 zsm_Latn Standard Malay 749 zul_Latn Zulu 838
@@ A.5 Analysis of ASR Performance
Language Whisper-v3-Large SeamlessM4Tv2-Large afr_Latn 13.1 12.6 amh_Ethi 207.8 23.9 arb_Arab 7.6 8.0 asm_Beng 99.0 15.1 ast_Latn 15.5 19.8 azj_Latn 7.2 8.3 bel_Cyrl 11.9 6.9 ben_Beng 33.9 9.0 bos_Latn 4.6 7.8 bul_Cyrl 5.2 9.6 cat_Latn 3.3 5.3 ceb_Latn 17.4 19.0 ces_Latn 4.2 8.0 ckb_Arab 98.3 13.9 cym_Latn 15.8 14.4 dan_Latn 4.8 10.0 deu_Latn 2.2 6.5 ell_Grek 6.4 10.8 eng_Latn 3.1 6.2 est_Latn 7.3 11.0 fin_Latn 3.2 10.2 fra_Latn 2.8 6.6 fuv_Latn 194.3 57.0 gaz_Latn 35.8 47.8 gle_Latn 85.1 21.4 glg_Latn 4.3 6.5 guj_Gujr 21.3 10.3 hau_Latn 33.8 51.8 heb_Hebr 12.2 14.5 hin_Deva 11.7 10.2 hrv_Latn 4.5 8.5 hun_Latn 5.2 8.5 hye_Armn 18.0 9.1 ibo_Latn 42.9 59.3 ind_Latn 4.0 9.2 isl_Latn 16.0 10.3 ita_Latn 2.1 5.2 jav_Latn 26.3 11.3 jpn_Jpan 6.3 16.4 kam_Latn 45.1 53.9 kan_Knda 20.1 10.9 kat_Geor 19.8 7.1 kaz_Cyrl 8.9 9.7 kea_Latn 35.8 38.1 khk_Cyrl 37.6 14.6 khm_Khmr 144.0 29.8 kir_Cyrl 28.3 8.8 kor_Hang 7.4 10.3 lao_Laoo 109.7 30.3 lin_Latn 22.0 58.7 lit_Latn 8.3 11.9
Language Whisper-v3-Large SeamlessM4Tv2-Large ltz_Latn 29.5 40.1 lug_Latn 44.9 14.1 luo_Latn 42.7 58.8 lvs_Latn 7.3 8.2 mal_Mlym 104.3 14.2 mar_Deva 24.3 10.7 mkd_Cyrl 5.8 8.6 mlt_Latn 26.0 12.1 mri_Latn 13.4 54.7 mya_Mymr 132.9 22.2 nld_Latn 3.1 6.8 nob_Latn 4.6 9.0 npi_Deva 25.6 74.1 nso_Latn 95.7 66.9 nya_Latn 35.6 12.4 oci_Latn 25.9 33.8 ory_Orya 93.8 11.4 pan_Guru 44.2 10.3 pbt_Arab 36.8 21.6 pes_Arab 8.6 7.7 pol_Latn 2.3 7.9 por_Latn 2.8 7.6 ron_Latn 3.2 6.6 rus_Cyrl 2.6 6.1 slk_Latn 3.9 7.0 slv_Latn 6.0 9.2 sna_Latn 26.7 32.7 snd_Arab 94.5 33.5 som_Latn 34.9 17.9 spa_Latn 2.2 6.3 srp_Cyrl 76.6 7.6 swe_Latn 3.7 9.2 swh_Latn 11.0 8.9 tam_Taml 13.3 11.6 tel_Telu 90.1 13.3 tgk_Cyrl 28.8 8.8 tgl_Latn 5.0 12.0 tha_Thai 9.8 11.4 tur_Latn 6.2 7.5 ukr_Cyrl 3.0 8.9 umb_Latn 156.7 55.5 urd_Arab 9.4 9.6 uzn_Latn 25.0 8.4 vie_Latn 4.7 6.7 wol_Latn 140.6 47.9 xho_Latn 48.1 32.3 yor_Latn 49.3 33.2 zho_Hans 16.2 19.0 zho_Hant 30.6 35.1 zsm_Latn 3.4 8.5 zul_Latn 42.6 18.4